Find the eccentricity of the ellipse 3x2 + 4y2 = 12 and the equation of the tangent to the ellipse at the point (1, 3/2). If this tangent meets the y-axis at the point G, and S and S′ are the foci of the ellipse, find the area of triangle SS′G.


1
Expert's answer
2021-08-17T10:36:02-0400
"3x^2+4y^2=12"

Differentiate both sides with respect to "x"


"6x+8yy'=0"

"y'=-\dfrac{3x}{4y}"

Point "(1, 3\/2)"


"y'(1, \dfrac{3}{2})=-\dfrac{3(1)}{4(\dfrac{3}{2})}=-\dfrac{1}{2}"

The equation of the tangent in point-slope form


"y-\dfrac{3}{2}=-\dfrac{1}{2}(x-1)"

The equation of the tangent in slope-intercept form


"y=-\dfrac{1}{2}x+2"


"3x^2+4y^2=12"

"\dfrac{x^2}{4}+\dfrac{y^2}{3}=1"

"a^2=4, b^2=3, c=\sqrt{a^2-b^2}=\sqrt{4-3}=1"



The eccentricity of the ellipse


"e=\dfrac{c}{a}=\dfrac{1}{2}"



Foci: "(\pm c, 0)"


"S(-1, 0), S'(1, 0)"

"x=0: y(0)=-\dfrac{1}{2}(0)+2=2"

"G(0,2)"




Triangle "SGS'"


"Area= \dfrac{a}{b}(2)(1-(-1))=2 (units^2)"

The area of triangle SS′G is 2 square units.